# GRASP: Replace Redundant Layers with Adaptive Singular Parameters for Efficient Model Compression

### v3

@@ Abstract
Recent studies have demonstrated that many layers are functionally redundant in large language models (LLMs), enabling model compression by removing these layers to reduce inference cost. While such approaches can improve efficiency, indiscriminate layer pruning often results in significant performance degradation. In this paper, we propose GRASP (Gradient-based Retention of Adaptive Singular Parameters), a novel compression framework that mitigates this issue by preserving sensitivity-aware singular values. Unlike direct layer pruning, GRASP leverages gradient-based attribution on a small calibration dataset to adaptively identify and retain critical singular components. By replacing redundant layers with only a minimal set of parameters, GRASP achieves efficient compression while maintaining strong performance with minimal overhead. Experiments across multiple LLMs show that GRASP consistently outperforms existing compression methods, achieving 90% of the original model’s performance under 20% compression ratio. The source code is available at GRASP.
@@ 2.2 Layer Replacement with Adaptive Singular Parameters
To address this limitation, GRASP introduces gradient-based attribution to evaluate the importance of each singular value based on its contribution to model performance, rather than relying on magnitude alone. Formally, the singular value decomposition of a weight matrix $W\in\mathbb{R}^{m\times n}$ is given by $W=U\Sigma V^{\top}$. Equivalently, $W$ can be represented as the sum of its rank-one components:
@@
where $T(\cdot)$ denotes the estimated change in loss when a certain parameter $\theta$ is zeroed out. This can be approximated by the second-order Taylor series expansion (LeCun et al., 1989), which is defined as:
@@
where $\nabla_{\theta}\mathcal{L}$ denotes the gradient of the standard language modeling objective function $\mathcal{L}$ with respect to the parameter $\theta$, and $H$ represents the corresponding Hessian matrix. To reduce computational overhead, we omit the second-order term and approximate the importance using only the first-order derivative (Hua et al., 2025, Kim et al., 2024):
@@
The first term in Eq. 6 captures the gradient projection of the loss $\mathcal{L}$ with respect to the singular value $\sigma_{i}$. The detailed proof is provided in Appendix A.1. By aggregating the first-order expansions across all components within each singular group, our method effectively captures the contribution of singular values. A larger value indicates a greater importance of the singular group. This gradient-oriented attribution moves beyond heuristic magnitude-based criteria, enabling performance-aligned importance evaluation. Under the hypothesis that redundant layers exhibit an inherent low-rank structure, we retain only the top-$k$ singular values most critical to model performance based on Eq. 6, and use them to replace the corresponding redundant layer.
@@ 2.3 Detailed Implementation of GRASP
4:Compute cosine similarity $\text{cos}(H_{i},H_{i+1})$ for all layers
@@
10:         Compute importance $I(\Phi_{k})$ for each $\Phi_{k}=\{u_{k},\sigma_{k},v_{k}^{\top}\}$ via Eq. 6
@@ 3.1 Experimental Setup
Table 2: Comparison between GRASP and structured pruning baselines with post-training compensation under a 20% compression ratio. Results marked with * are reported from (Chen et al., 2024). Bold values indicate the best performance.
@@ 3.2 Comparison with Pruning-based LLM Compression Methods
3.2.1 Experiment 1
@@
3.2.2 Experiment 2
@@
3.2.3 Experiment 3
@@ 3.3 Compression Costs and Inference Efficiency of GRASP
GRASP enables low-cost compression of LLMs while improving inference efficiency on real hardware. To evaluate its acceleration benefits, we measure the throughput (tokens per second) of the original LLaMA2-7B and its GRASP-compressed counterpart under varying sequence lengths and batch sizes. As shown in Figure 5, GRASP consistently improves generation speed and achieves acceleration comparable to direct layer removal. Notably, although GRASP retains a small subset of parameters within redundant layers to mitigate the performance drop caused by layer removal, these retained components are extremely low-rank and incur negligible inference overhead while preserving task performance. Additionally, we measure the compression time of GRASP and other structured pruning baselines when compressing LLaMA2-7B on an NVIDIA A100 GPU under a 25% compression ratio. As shown in Table 4, GRASP is able to compress the model efficiently while maintaining strong downstream performance, demonstrating its practicality for real-world deployment.
@@ 3.4 Ablation Study
Table 5: Comparison of GRASP using different types and amounts of data for compressing LLaMA3.1-8B-Instruct at a 20% compression ratio. Results are reported on the WikiText-2, PTB datasets (measured by $\downarrow$) and the average accuracy across seven commonsense reasoning benchmarks: OpenBookQA, WinoGrande, HellaSwag, ARC-easy, ARC-challenge, PIQA, and MathQA.
@@
As detailed in Section 2.3, GRASP processes the redundant layers in a sequential manner, starting from the final redundant layer and proceeding backward through the network. This can be done using one-shot pruning, which identifies and decomposes all redundant layers in a single step, or iterative pruning, which processes layers one at a time to account for interactions between layers. We present the ablation results in Table 10 in Appendix A.8, which shows that both approaches achieve similar performance, with one-shot pruning being more efficient.
@@ 5 Related Work
Another line of structured pruning research focuses on low-rank approximation, where Singular Value Decomposition (SVD) is widely used in model compression to decompose weight matrices into low-rank structures, typically selecting top-k singular values based on Frobenius norm reconstruction loss. Recent methods have enhanced SVD to reduce LLM compression error. FWSVD (Hsu et al., 2022) incorporates Fisher information to reweight the importance of parameters before applying SVD. ASVD (Yuan et al., 2023) uses activation patterns from a calibration dataset to scale weight matrices, reducing compression-induced activation errors. SVD-LLM (Wang et al., 2024) applies truncation-aware data whitening and layer-wise updates to ensure a direct relationship between singular values and compression loss. Additionally, (Yu and Wu, 2023; Chavan et al., 2024; Ji et al., 2024) present another paradigm for low-rank compression of LLMs, where eigenvalue decomposition is applied to output activations, approximating the output features of each layer with low-rank matrices. In contrast to these approaches, GRASP integrates gradient-based attribution into the low-rank decomposition process, enabling both efficient compression and critical parameter retention.
@@ A.2 Experimental Setup and Hyperparameters Configuration
A.3 Comparison of Concurrent Structured Pruning Methods
@@
A.4 Detailed Results on Commonsense Reasoning Benchmarks
@@
A.5 More Results on Other Models
@@
A.6 Evaluation Results on LongBench
@@
A.7 Robustness of GRASP towards different calibration Dataset
In this section, we provide details of the ablation studies conducted to investigate the impact of calibration datasets and the amount of data used for singular value gradient attribution. Specifically, we selected 512 samples from WikiText-2 (Merity et al., 2017) and C4 (Raffel et al., 2020) as calibration data to assess the performance of GRASP when compressing LLaMA3.1-8B-Instruct under 20% compression ratio. Additionally, we selected 64, 128, 256 and 512 samples from WikiText-2 to examine the robustness of GRASP to the change in the number of calibration data. All calibration data were randomly selected from the training splits of the downstream datasets, ensuring no data leakage. As shown in Figure 4, we can observe that GRASP consistently achieves strong performance, indicating that our method is robust to variations in both the calibration dataset and the number of data. Tables 12 and 13 summarize the results of GRASP when compressing LLaMA3.1-8B-Instruct with different calibration datasets (WikiText-2, C4) and varying numbers of calibration data.
A.8 Robustness of GRASP towards Different Pruning Strategy
In this section, we present the results of ablation experiments conducted with different pruning strategies in GRASP, with the results summarized in Table 10
Pruning Strategy WikiText-2 PTB
Average
Accuracy
Compression
Time(h)
One-shot Pruning 37.86 63.97 47.12 0.16
Iterative Pruning 38.39 72.18 47.13 0.22

### v4

@@ Abstract
Recent studies have demonstrated that many layers are functionally redundant in large language models (LLMs), enabling model compression by removing these layers to reduce inference cost. While such approaches can improve efficiency, indiscriminate layer pruning often results in significant performance degradation. In this paper, we propose GRASP (Gradient-based Retention of Adaptive Singular Parameters), a novel compression framework that mitigates this issue by preserving sensitivity-aware singular values. Unlike direct layer pruning, GRASP leverages gradient-based attribution on a small calibration dataset to adaptively identify and retain critical singular components. By replacing redundant layers with only a minimal set of parameters, GRASP achieves efficient compression while maintaining strong performance with minimal overhead. Experiments across multiple LLMs show that GRASP consistently outperforms existing compression methods, achieving 90% of the original model’s performance under 20% compression ratio. The source code is available at https://github.com/LyoAI/GRASP.
@@ 2.2 Layer Replacement with Adaptive Singular Parameters
To address this limitation, GRASP introduces gradient-based attribution to evaluate the importance of each singular value by its contribution to model performance, rather than relying on magnitude alone. Formally, the singular value decomposition of a weight matrix $W\in\mathbb{R}^{m\times n}$ is given by $W=U\Sigma V^{\top}$. Equivalently, $W$ can be expressed as the sum of its rank-one components:
@@
where $T(\cdot)$ denotes the estimated loss change when a certain parameter $\theta$ is zeroed out. This can be approximated by the second-order Taylor expansion (LeCun et al., 1989), defined as:
@@
where $\nabla_{\theta}\mathcal{L}$ denotes the gradient of the standard language modeling objective function $\mathcal{L}$ w.r.t. the parameter $\theta$, and $H$ represents the corresponding Hessian matrix. To reduce computational overhead, we omit the second-order term and approximate the importance using only the first-order derivative (Hua et al., 2025, Kim et al., 2024):
@@
The first term in Eq. 6 captures the gradient projection of the loss $\mathcal{L}$ w.r.t. the singular value $\sigma_{i}$. The proof is given in AppendixA.1. By aggregating the first-order expansions across all components within each singular group, our method effectively captures the contribution of singular values, where larger values indicate greater importance. This gradient-oriented attribution moves beyond heuristic magnitude-based criteria, enabling performance-aligned importance evaluation. Under the hypothesis that redundant layers exhibit an inherent low-rank structure, we retain only the top-rr% singular values most critical to model performance based on Eq. 6, and use them to replace the corresponding redundant layer. The relationship between the singular group retain ratio (rr%) and the overall target compression ratio is detailed in Appendix A.3.
@@ 2.3 Detailed Implementation of GRASP
4:Compute cosine similarity $\text{cos}(H_{i},H_{i+1})$ for all layers via Eq.1 using $D$
@@
10:   Compute importance $I(\Phi_{k})$ for each $\Phi_{k}=\{u_{k},\sigma_{k},v_{k}^{\top}\}$ via Eq.6 using $D$
@@ 3.1 Experimental Setup
Table 2: Comparison between GRASP and structured pruning baselines with post-training compensation under a 25% compression ratio. Results marked with * are reported from (Chen et al., 2024). Bold values indicate the best performance.
@@ 3.2 Comparison with Pruning-based LLM Compression Methods
3.2.1 Comparison without Post-Training Compensation
@@
3.2.2 Comparison with Post-Training Compensation
@@
3.2.3 Comparison with Low-Rank Pruning Methods
@@ 3.3 Compression Costs and Inference Efficiency of GRASP
GRASP enables low-cost compression of LLMs while improving inference efficiency on real hardware. To evaluate its acceleration benefits, we measure the throughput (tokens per second) of the original LLaMA2-7B and its GRASP-compressed counterpart under varying sequence lengths and batch sizes. As shown in Figure 5, GRASP consistently improves generation speed and achieves acceleration comparable to direct layer removal. Notably, although GRASP retains a small subset of parameters within redundant layers to mitigate the performance drop caused by layer removal, these retained components are extremely low-rank and incur negligible inference overhead while preserving task performance. Additionally, we measure the compression time of GRASP and other structured pruning baselines when compressing LLaMA2-7B on an NVIDIA A100 GPU under a 25% compression ratio. As reported in Table 4, Pruning Time refers to the time required to perform model compression up to the target sparsity, excluding any compensation. Compensation Time refers to the time spent on post-compression procedures—such as fine-tuning or parameter updates—aimed at recovering model performance. The results show that GRASP is able to compress the model efficiently while maintaining strong performance, demonstrating its practicality for real-world deployment.
@@ 3.4 Ablation Study
[…]
Retain Ratio rr%.
We further explore the impact of the retain ratio $r$ on GRASP’s performance. As shown in Table 6 and Figure 6, performance improves rapidly when increasing $r$ from 0% to 10%, after which the gains largely saturate. This indicates that a retain ratio of 10% serves as a practical inflection point, providing a strong trade-off between compression and accuracy. In practice, we therefore adopt a fixed retain ratio of 10%, which consistently yields stable performance across benchmarks while maintaining high compression efficiency. These results further confirm that the selected redundant layers contain substantial redundancy, as even a small fraction of retained singular components suffices to recover most of the original performance.
Retain Ratio
Overall
Compression Ratio
Openb. ARC_e ARC_c WinoGrande PIQA MathQA Average
0% 24.0% 20.2 45.1 30.9 61.3 67.0 22.2 41.1
5% 22.8% 22.4 54.9 31.8 66.0 70.0 23.8 44.8
10% 21.6% 23.2 56.1 32.7 66.2 70.7 23.9 45.5
15% 20.4% 23.4 56.3 32.9 65.6 70.7 24.1 45.5
20% 19.2% 23.4 56.5 33.4 65.6 71.5 24.3 45.8
Table 6: Ablation results on LLaMA2-7B by replacing 8 redundant layers with sensitivity-aware singular parameters while varying the retain ratio. The table reports the resulting overall compression ratio and downstream task performance across multiple benchmarks.
Figure 6: Effect of the retain ratio on average performance when compressing 8 redundant layers of LLaMA2-7B. The results show a sharp improvement up to a 10% retain ratio, after which performance gains plateau, indicating 10% as a practical inflection point.
@@
As detailed in Section 2.3, GRASP processes the redundant layers in a sequential manner, starting from the final redundant layer and proceeding backward through the network. This can be done using one-shot pruning, which identifies and decomposes all redundant layers in a single step, or iterative pruning, which processes layers one at a time to account for interactions between layers. We present the ablation results in Table 7, which shows that both approaches achieve similar performance, with one-shot pruning being more efficient.
Pruning Strategy WikiText-2 PTB
Average
Accuracy
Compression
Time(h)
One-shot Pruning 37.86 63.97 47.12 0.16
Iterative Pruning 38.39 72.18 47.13 0.22
@@ 5 Related Work
Another line of structured pruning research focuses on low-rank approximation, where Singular Value Decomposition (SVD) is widely used to decompose weight matrices into low-rank structures, typically selecting top-k singular values based on Frobenius norm reconstruction loss. Recent methods have enhanced SVD to reduce LLM compression error. FWSVD (Hsu et al., 2022) incorporates Fisher information to reweight the importance of parameters before applying SVD. ASVD (Yuan et al., 2023) uses activation patterns from a calibration dataset to scale weight matrices, reducing compression-induced activation errors. SVD-LLM (Wang et al., 2024) applies truncation-aware data whitening and layer-wise updates to ensure a direct relationship between singular values and compression loss. Additionally, (Yu and Wu, 2023; Chavan et al., 2024; Ji et al., 2024) present another paradigm for low-rank compression of LLMs, where eigenvalue decomposition is applied to output activations, approximating the output activations with low-rank matrices. However, SVD-based methods require truncating at least 50% of singular values to reduce parameters of square matrices (which are common in LLMs like Llama), which often leads to significant information loss. In contrast to these approaches, GRASP integrates gradient-based attribution into the low-rank decomposition process and replaces redundant layers with only a small fraction of parameters (typically 10%), thereby enabling efficient compression while retaining critical information.
@@ A.2 Experimental Setup and Hyperparameters Configuration
A.3 Relationship between Singular Group Retail Ratio and Target Overall Compression Ratio
In GRASP, the retain ratio rr% denotes the proportion of singular components preserved within each weight matrix of the selected redundant layers. The target overall model compression ratio, which is jointly determined by (i) the number of layers selected in Step 1, and (ii) the retain ratio rr% applied in Step 2.
In practice, we adopt a fixed retain ratio—typically 10%—as ablation studies (Section 3.4) show it achieves a favorable trade-off between compression and performance. Table 6 further illustrates how varying the retain ratio influences the overall compression ratio and downstream performance across several benchmarks when compressing 8 layers of LLaMA2-7B. These results highlight the high redundancy of the selected layers: even a small fraction of retained singular components is sufficient to recover most of the original performance.
A.4 Comparison of Concurrent Structured Pruning Methods
@@
A.5 Detailed Results on Commonsense Reasoning Benchmarks
@@
A.6 More Results on Other Models
@@
A.7 Evaluation Results on LongBench
@@
A.8 Robustness of GRASP towards different calibration Dataset
In this section, we provide details of the ablation studies conducted to investigate the impact of calibration datasets and the amount of data used for singular value gradient attribution. Specifically, we selected 512 samples from WikiText-2 (Merity et al., 2017) and C4 (Raffel et al., 2020) as calibration data to assess the performance of GRASP when compressing LLaMA3.1-8B-Instruct under 20% compression ratio. Additionally, we selected 64, 128, 256 and 512 samples from WikiText-2 to examine the robustness of GRASP to the change in the number of calibration data. All calibration data were randomly selected from the training splits of the downstream datasets, ensuring no data leakage. As shown in Figure 4, we can observe that GRASP consistently achieves strong performance, indicating that our method is robust to variations in both the calibration dataset and the number of data. Tables 13 and 14 summarize the results of GRASP when compressing LLaMA3.1-8B-Instruct with different calibration datasets (WikiText-2, C4) and varying numbers of calibration data.
A.9 Comparison with Unstructured Pruning Methods
To provide a more comprehensive picture of model compression, we have included comparisons with two widely used unstructured pruning methods, SparseGPT (Frantar and Alistarh, 2023) and Wanda (Sun et al., 2023) in our experiments. While unstructured pruning introduces fine-grained sparsity in weight matrices, it typically requires dedicated sparse formats to realize storage or runtime benefits, which complicates direct comparison under the same compression ratio. To ensure a fair evaluation, we adopt the commonly used 2:4 sparsity pattern (i.e., 50% sparsity) for both methods, and compare them against GRASP at a 20% structured compression ratio. The results are summarized in Table 15, indicating that GRASP achieves competitive or superior performance against unstructured pruning methods across all benchmarks.
@@ A.9 Comparison with Unstructured Pruning Methods
Method Openb. ARC_e ARC_c WinoG. HeSW PIQA MathQA Average
Dense 32.12 72.69 39.87 67.33 57.16 78.46 26.02 53.38
SparseGPT(2:4) 24.40 64.10 30.80 66.61 43.09 70.73 24.25 46.28
wanda(2:4) 24.60 62.75 30.80 62.67 41.38 70.29 23.62 45.16
GRASP 24.80 56.02 32.68 66.14 42.57 70.74 23.92 45.27
GRASP* 28.20 68.73 37.63 67.88 50.98 72.47 24.19 50.01
Table 15: Comparison between GRASP and unstructured pruning methods under the 2:4 sparsity pattern (50% sparsity). GRASP is applied at a 20% structured compression ratio. GRASP* denotes the variant with light performance compensation